OpenStudy (anonymous):

for what value(s) of (a) the function a^x and y=1+x tangent at x=0

OpenStudy (inkyvoyd):

Uhm, I don't understand - can you reword that please?

OpenStudy (anonymous):

when is y=A^x tangent to Y=1+x at x=0

OpenStudy (anonymous):

the derivativ of a^x is a^x ln(a) at zero is ln(a) Now ln(a) must be equal to the slope of y=x+1 so \[ \ln(a) =1\\ a=e \]

OpenStudy (anonymous):

oh, duh.... (slap myself on the head)... good job @eliassaab .